A candidate was to substract 15 from a certain number, but mistakely added 25 and his answer was 145 find the percentage error

Answer by Edwin McCravy(20063)


Let the certain number be x

What he should have gotten = x-15

What he actually got = x+25, which is given = 145

So:

          x+25 = 145
             x = 120

Therefore:

What he should have gotten = x-15 = 120-15 = 105
   




Edwin